Derive the expression for impedance which varies along the transmission line. A 30-m-long lossless transmission line with Zo = 50  operating at 2 MHz is terminated with a load ZL =70+j30. Find the reflection coefficient and standing wave ratio.


1
Expert's answer
2021-04-26T17:00:57-0400

Answer

Zo = 50 

ZL =70+j30

So reflection coefficient

"\tau=\frac{Z_L-Z_0}{Z_L+Z_0}\\=\frac{ 20+j30 }{ 120+j30 }"

And also

Standing wave ratio

"SWR=\frac{\tau+1}{\tau-1}\\=\frac{(\frac{ 20+j30 }{ 120+j30 }+1)}{(\frac{ 20+j30 }{ 120+j30 }-1)}\\"

"=\frac{20+j30+120+j30}{20+j30-120-j30}"

"\\=

\frac{140+j60}{-100}"